company buys machinery for USD 60,000 on December 31, Year 1. Depreciation using the straight line method is to be applied. It is estimated that the machine's useful life will be six years and its residual value will be USD 6,000. Two years later (December 31, Year 3), the useful life is revised downward to three remaining years. The machine is sold for USD 25,000 on December 31, Year 5. What is the profit on disposal?
Answer #1

Depreciation = Value of Asset - Salvage Value / Estimated Useful Life in Years

Value of Machinery on Dec 31 Year 1 = $ 60,000

Salvage Value = $ 6,000

Estimated Useful Life in Years = 6 Years

Depreciation = 60,000 - 6,000 / 6

Depreciation = 54,000 / 6,000

Depreciation = $ 9,000 per Year

Value of Machinery on Dec 31, Year 3 = Value of Machinery on Dec 31, Year 1 - Depreciation for Year 2 - Depreciation for Year 3

Value of Machinery as on Dec 31 Year 3 = 60,000 - 9,000 - 9,000

Value of Machinery as on Dec 31 Year 3 = $ 42,000

Change in Estimates on Dec 31 Year 3

Value of Machinery as on Dec 31 Year 3 = $ 42,000

Salvage Value = $ 6,000

Remaining Useful Life in Years = 3 Year ( As per Change in Estimates)

Revised Depreciation Expense = 42,000 - 6,000 / 3

Revised Depreciation Expense = 36,000 / 3

Revised Depreciation Expense = $ 12,000 per Year for remaining 3 Years

Value of Machinery as on Dec 31 Year 5 = Value of Machinery as on Dec 31 Year 3 - Depreciation Expense for Year 4 - Depreciation Expense for Year 5

Value of Machinery as on Dec 31 Year 5 = 42,000 - 12,000 - 12,000

Value of Machinery as on Dec 31 Year 5 = $ 18,000

Profit on Disposal of Machinery = Sale Price of Machinery - Value of Machinery as on Dec 31 Year 5

Sale Price of Machinery = $ 25,000

Profit on Disposal of Machinery = 25,000 - 18,000

Profit on Disposal of Machinery = $ 7,000